APT and Dpkg Quick Reference Sheet  Matthew Danish

Common APT usage

apt-get install ⟨package⟩ Downloads ⟨package⟩ and all of its dependencies, and installs or upgrades them. This will also take a package off of hold if it was put on. See below for more info on hold.

apt-get remove [--purge] ⟨package⟩ Removes ⟨package⟩ and any packages that depend on it. -- purge specifies that packages should be purged, see dpkg -P for more information.

apt-get update Updates packages listings from Debian mirrors, should be run at least once a day if you install anything that day, and every time after /etc/apt/sources.list is changed.

apt-get upgrade [-u] Upgrades all packages installed to newest versions available. Will not install new or remove old packages. If a package changes dependencies and requires installation of a new package, it will not be upgraded, it will be put on hold instead. apt-get upgrade will not upgrade packages put on hold (that is the meaning of hold). See below for how to manually put packages on hold. I suggest the ‘-u’ option as well, because then you can see what packages are going to be upgraded.

apt-get dist-upgrade [-u] Similar to apt-get upgrade, except that dist-upgrade will install or remove packages to satisfy dependencies.

apt-cache search ⟨pattern⟩ Searches packages and descriptions for ⟨pattern⟩.

apt-cache show ⟨package⟩ Shows the full description of ⟨package⟩.

apt-cache showpkg ⟨package⟩ Shows a lot more detail about ⟨package⟩, and its relationships to other packages. synaptic gdebi gdebi-kde dselect aptitude gnome-apt Graphical front ends to APT (some of these may be in their own package, that must be installed before use). While dselect is arguably the most powerful, it’s also the oldest and hardest to use.

Common Dpkg usage

dpkg -i ⟨package.deb⟩ Installs a Debian package file; one that you downloaded manually, for example.

dpkg -c ⟨package.deb⟩ Lists the contents of ⟨package.deb⟩, a .deb file.

dpkg -I ⟨package.deb⟩ Extracts package information from ⟨package.deb⟩, a .deb file.

dpkg -r ⟨package⟩ Removes an installed package named ⟨package⟩

dpkg -P ⟨package⟩ Purges an installed package named ⟨package⟩. The difference between remove and purge is that while remove only deletes data and executables, purge also deletes all configuration files in addition.

dpkg -L ⟨package⟩ Gives a listing of all the files installed by ⟨package⟩. See also dpkg -c for checking the contents of a .deb file.

dpkg -s ⟨package⟩ Shows information on the installed package ⟨package⟩. See also apt-cache show for viewing package information in the Debian archive and dpkg -I for viewing package information extracted from a .deb file.

dpkg-reconfigure ⟨package⟩ Reconfigures an installed package, if it uses debconf (debconf provides that consistent configuration interface for package installation). You can reconfigure debconf itself if you want to change the front-end or priority of questions asked. For example, to reconfigure debconf with the dialog front-end, you simply run:

dpkg-reconfigure --frontend=dialog debconf

echo “⟨package⟩ hold” | dpkg --set-selections Put ⟨package⟩ on hold (command line method)

dpkg --get-selections “⟨package⟩” Get the current status of ⟨package⟩ (command line method)

dpkg -S ⟨file⟩ Searches for ⟨file⟩ in package database, telling you which installed packages have that file in them. To search for a file or list the contents of a package without installing it, you can use apt-file command line tool.

Building Debian packages from Source

apt-get source [-b] ⟨package⟩ Download the source Debian package for ⟨package⟩ and extract it. You must have deb-src lines in your /etc/apt/sources.list for this to work. If you supply the ‘-b’ option and you are currently root, then the package will be automatically built if possible.

apt-get build-dep ⟨package⟩ Download and install the packages necessary to build the source Debian package ⟨package⟩. This feature is only present in apt version 0.5 and up. Currently this means that woody and above contain this functionality. If you have an older version of apt then the easiest way to find out the build dependencies is to look in the debian/control file in the source package directory. A common usage of this command is in conjunction with apt-get source -b. For example (as root):

apt-get build-dep ⟨package⟩ apt-get source -b ⟨package⟩

Will download the source package, all of its build dependencies, and attempt to compile the source package. dpkg-source -x ⟨package.dsc⟩ If you have downloaded the source package for a program manually, which includes several files such as a .orig.tar.gz (or .tar.gz if it is Debian native), a .dsc, and a .diff.gz (if it is not Debian native), then you can unpack the source package using this command on the .dsc file.

dpkg-buildpackage Builds a Debian package from a Debian source tree. You must be in the main directory of the source tree for this to work. Sample usage:

dpkg-buildpackage -rfakeroot -uc -b

Where ‘-rfakeroot’ instructs it to use the fakeroot program to simulate root privileges (for ownership purposes), ‘-uc’ stands for “Don’t cryptographically sign the changelog”, and ‘-b’ stands for “Build the binary package only”

debuild A handy wrapper script around dpkg-buildpackage that will automatically take care of using fakeroot or not, as well as running lintian and gpg for you. This script is provided by (:program ), so you need to install this package first.

Fixing dependencies

dpkg --configure --pending If dpkg quits with

an error while apt-get install, upgrade, or

dist-upgradeing try running this to configure the packages that were already unpacked. Then try

apt-get install, upgrade, or dist-upgrade
-f, and then try apt-get install, upgrade,

or dist-upgrade again. Repeat as needed. This usually resolves most dependency problems (also, if it mentions a specific package for some reason, you might want to try installing or removing that package)

apt-get install -f apt-get upgrade -f apt-get dist-upgrade -f Attempt to fix dependencies while doing one of the above. Note that apt-get install -f does not require a ⟨package⟩ argument.

See Also

man ⟨package⟩ To learn more about this commands you can always consult their man pages, e.g. man apt-get, man dpkg. You can also find information about this and more tools, HOWTOs, manuals, etc at http://www.debian.org/doc/.